In my class, one-half of the students went to the dance, one-third of the remainder went to the movies, and the remaining 8 people stayed home. How many students are in my class?
Answer by checkley77(12569)

X/2+1/3*X/2+8=X
X/2+X/6-X+8=0
(3X+X-6X)/6=-8
-2X=-48
X=-48/-2
X=24 TOTAL NUMBER OF STUDENTS.
PROOF:
24/2+1/3*24/2+8=24
12+24/6+8=24
12+4+8=24
24=24